OpenStudy (anonymous):

Joanne works two jobs to pay for college. She tutors for $25 per hour and also works as a receptionist for $12 per hour. Due to her class and study schedule, Joanne is only able to work up to 25 hours per week, but must earn at least $200 per week. If t represents the number of hours Joanne tutors and r represents the number of hours she works as a receptionist, which system of inequalities represents this scenario?

OpenStudy (anonymous):

t + r greater than or equal to 25 25t + 12r = 200 t + r less than or greater to 25 25t + 12r less than or greater to 200 t + r less than or greater to 25 25t + 12r greater than or equal to 200 None of the systems shown represent this scenario.

OpenStudy (anonymous):

For sure, it should be t+r<=25

OpenStudy (anonymous):

It is the 25t+12r greater than or = to 200 because t=hours as tutor and r=hours as receptionist. She makes $25 per hour as tutor, 25t and $12 per hour as receptionist, 12r. Together (the sum) she needs AT LEAST 200 so the sum of 25t and 12r must be greater than or equal to 200

OpenStudy (anonymous):

Thanks soo much! :)

OpenStudy (anonymous):

The first part should be t+r is less than or equal to 25 though since she has to work up to 25 hours

OpenStudy (anonymous):

Thanks for that!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

=^.^=